# Unlocking Text Capabilities in Vision Models

### v2

@@ 5 Experiments
(b) Zero-Shot Compositional Captioning Performance
Table 3: Comparison of Zero-Shot Captioning Performance
Figure 5: Qualitative examples of decoded visual features from different visual classifiers, into text.
Note that our results in Table 3(a) are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics. However, it is important to note that B4 and M are n-gram overlap-based metrics. They assume that the generated caption follows a specific structure and style. We verify this by revisiting an old image captioning paradigm termed as compositional captioning [52], and later revived with deep learning methods [53]. In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence. With the current advancements of Large Language Models (LLMs), we can use an LLM as a compositioner. Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 4.1 and shown in Figure 3(a), and feed them, along with their similarity scores, to an LLM. For the concepts, we use the same concept set as in Section 4.1. We also add a list of the most common verbs [54] in English to the pool. This allows us to cover all possible words and interactions. We prompt the LLM to utilize the provided information to compose a sentence, given a very few in-context examples from the COCO captioning training set (in our experiments, we use 9 examples). This allows us to generate sentences adhering to a specific style and structure. For this experiment, we used GPT4o-mini [55] as our LLM, as it is fast and cost-efficient. In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importances. We provide details on the exact prompt we used in Appendix Section J. Results are shown in Table 3(b). While results on CiDEr and SPICE are incremental compared to results in Table 3(a), the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M in Table 3(a) compared to baseline methods. Also note that, since the LLM is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains (see Section J.1 in the appendix).
6 Conclusion
We presented a method to transform visual classifiers such that they can be queried via open-set text queries. We achieve state-of-the-art results in two applications, i.e., zero-shot CBMs and decoding of visual features into natural language sentences. Our method can be applied to various other applications that were previously restricted to CLIP models. Our work removes this restriction, being applicable to any visual classifier. Finally, as with any research work, this study has its own limitations which are discussed in Appendix Section A.
Appendix A Limitations
As with any research work, our study has its own limitations that should be transparent and acknowledged. In particular, we identify two primary limitations of our method: the first concerns wrong semantic associations of class names in CBMs, while the second pertains to the limited generalization of our method to fine-grained datasets.
We start by addressing the first limitation of wrong semantic associations of class names in CBMs. This issue is closely tied to the choice of concept set used. When using the 20K most common words in English as our concept set, we observe that class names get associated to wrong semantically-related concepts. Figure 1 illustrates such cases. In the first example, the bird drake is linked to artist-related concepts (Rihanna, Robbie, lyric). This occurs because the bird “drake" is less familiar to the text encoder than the artist “drake". In fact, a google search with the word “drake” yields directly the artist rather than the bird. In the second example, the top-detected concepts for the prediction “african grey” are incorrect semantic associations with the word “african” (ethiopian, tanzania) and the word "grey" (purple, blue) that do not pertain to the bird itself. We also observe similar cases that may raise ethical concerns (e.g., the animal “cock" leads to associations with male reproductive terms). For each example, we also report the total logit score of the prediction.
However, it is worth noting the following:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Figure 1: Limitations of our method is wrong semantic concept association in CBMs
The second limitation is the lack of generalization to fine-grained datasets. We test our ImageNet-trained reformulated classifiers to perform zero-shot transfer to the Places365 dataset [56] for scene classification. It is important to note that Places365 is known to induce low performance, even when using supervised training with the powerful CLIP model, making it a challenging classification dataset. To perform zero-shot transfer, we convert the class names of Places365 into text prompts of the form: “a scene of a {class} location" which are then encoded using the text encoder $T$ to obtain the classification weights $U_{places}$. The image is encoded with the respective vision encoder to obtain the visual features $f$, which are then projected via the MLP to yield $\tilde{f}$. Zero-shot classification is performed using $\text{argmax}(\tilde{f}.U^{T}_{places})$. Results on the Places365 validation set are shown in Table 1. We report the supervised learning performance by training a linear probe on top of the CLIP model features. The supervised top-1 accuracy is 53.40 and 55.10 for ResNet50 and ViT-B/16 CLIP models, respectively. As demonstrated, these low top-1 accuracies suggest that Places365 is a particularly challenging dataset for classification. We then report the zero-shot performance of our method on several visual classifiers. The best-performing model achieves a top-1 zero-shot accuracy of 14.36%, much lower than the supervised performance.
However, we emphasize that the observed low generalization performance is not attributed to our formulation; rather, it stems from the inherent limitations of the classifier itself. If the classifier (e.g., ResNet101) fails to generalize to out-of-distribution (OOD) fine-grained classes—as is the case—then it is unreasonable to expect the text-transformed classifier to perform any better. Since our transformed classifier is a replication of the original classifier, it inherits all of its limitations. In fact, expecting improved performance from the transformed classifier would be counterintuitive, as it would indicate a departure from the original classifier’s underlying reasoning. Moreover, as demonstrated in Table 1, models that benefit from large-scale pretraining tend to achieve superior OOD fine-grained performance, a trend that is consistent with findings in the OOD literature [57]. Finally, our approach performs significantly better than a randomly shuffled classifier where we shuffle its weights across classes.
Model Top-1 (%) Top-5 (%)
Supervised
CLIP RN50 Linear Probe 53.40 -
CLIP ViT-B Linear Probe 55.10 -
Zero-Shot
Shuffled Classifier 0.24 1.54
DenseNet161 10.55 25.18
ResNet50 10.58 25.14
ResNet101 10.97 25.42
ViT-B/16 11.47 26.53
ResNet50v2 11.66 27.40
ConvNeXt-Base 11.76 27.12
Swin-Base 12.12 26.83
Swinv2-Base 12.13 27.18
EfficientNetv2-M 12.31 28.33
ViT-B/16v2 13.11 29.54
DINOv2-Base 13.96 32.35
ConvNeXtv2 (pt) 14.05 30.92
BeiT-L/16 14.15 31.08
ViT-B/16 (pt) 14.22 32.18
ConvNeXt (pt) 14.36 31.64
Table 1: Top-1 and Top-5 zero-shot transfer results to Places365 for various models
Appendix B Pseudocode of our method
We provide a PyTorch-like pseudocode of our training approach in Listing 1.
# mlp: trainable MLP from visual_dim -> text_dim
# images: batch of B images, shape (N, 3, height, width)
visual_feats = vision_encoder(images) # (N, visual_dim)
logits = visual_feats @ classifier # (N, num_classes)
original_dist = softmax(logits, dim=-1) # (N, num_classes)
mapped_feats = mlp(visual_feats) # (N, text_dim)
mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
text_feats = l2_norm(text_feats) # (N, text_dim)
pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
# cross entropy with original model’s soft distribution
loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
loss.backward() # only mapper parameters are updated
Listing 1: PyTorch-like pseudocode for our method
Appendix C Ablation Studies of the MLP
In this section, we evaluate the impact of the MLP and verify its effectiveness. We perform the following ablation studies: 1) Mean Ablation: For an image, we replace the input features to the MLP with a constant mean of the features calculated across the full ImageNet validation set. 2) Random Features: For an image, we replace the input features to the MLP with random values sampled from a normal distribution with a mean and standard deviation equal to that of the features calculated across the full ImageNet validation set. 3) Random Weight Ablation: We randomize the weights of the MLP projection. 4) Shuffled Ablation: For an image, we replace the input features to the MLP with input features of another random image in the validation dataset.
ResNet101v2 Ours 81.49 81.49 81.49 81.49
Ablated 0.10 0.11 1.70 0.11
ConvNeXt-Base Ours 83.88 83.88 83.88 83.88
Ablated 0.10 0.11 1.79 0.10
BeiT-L/16 Ours 87.22 87.22 87.22 87.22
Ablated 0.10 0.11 1.87 0.11
DINOv2-B Ours 84.40 84.40 84.40 84.40
Ablated 0.10 0.13 1.76 0.09
Table 2: Ablation studies of the MLP.
Appendix D Ablation Studies on the Text Encoder
In Table 3, we present ablation studies using other text encoders from the Sentence-BERT library [58] with the ResNet50 visual classifier. We observe that the choice of the text encoder has minimal effect on the performance. This is because even lower-performing text encoders are capable of understanding class names.
Text Encoder Top-1 (%)
DistilRoberta 75.73
MPNet-Base 75.78
MPNet-Base-MultiQA 75.76
MiniLM 75.80
Table 3: Ablation studies on other text encoders
Appendix E Effectiveness of Concept Interventions
As a supplementary experiment, we also report concept intervention results on our ZS-CBMs. Interventions on CBMs are an effective tool to mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions. The Waterbirds-100 dataset [59] is a standard dataset used in previous works [18] to conduct CBM intervention experiments. It is a binary classification dataset of two classes: waterbirds and landbirds. The training images of waterbirds are on water backgrounds, and training images of landbirds are on land backgrounds. However, the validation images do not have that correlation, where waterbird images appear on land backgrounds and landbird images on water backgrounds. The model is assumed to learn the water–land background correlation to perform this classification task. By building a CBM, we can correct this bias by intervening in concepts in the CB layer. However, there are two challenges associated with using the Waterbirds dataset in our work, and therefore we curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set. We provide more details about this process in Section L. Our curated dataset includes 140 validation images (70 for each class). We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from [18] which includes a collection of bird-related concepts and a collection of land-related concepts. The ZS-CBM achieves a low accuracy, as shown in Table 4, which indicates the water-land bias the model performs for classification. To correct this, we intervene in the concepts in the CB layer, following the setup from [18]. Intervention R (Int. R): We zero-out activations of any bird concepts from the bottleneck layer, and expect the accuracy to drop. Intervention K (Int. K): We keep activations of bird concepts as they are, but scale down the activations of all remaining concepts by multiplying them with a factor of 0.1, and expect the accuracy to increase. Results are presented in Table 4 for some models, and demonstrate the success of our intervention experiments.
We also conduct concept intervention experiments with a more challenging multi-class setup. Unlike the Waterbirds dataset where bias-correlation issues are assumed, we make no such assumption here. We instead evaluate whether intervening on class-related concepts in the bottleneck layer and zeroing out their activations, impairs model accuracy. A drop in accuracy indicates that these concepts are important for prediction. We select a subset of 10 classes from ImageNet following [60]. We use the original ImageNet validation images for those classes, rather than the validation set from [60], because the latter contains original ImageNet training examples that our model has already seen. We employ an LLM to generate 5 highly-relevant concepts for each class, achieving in total 50 concepts. The classes and concepts we used are provided in Section M of the supplementary material. We then measure CBM accuracy before and after intervention. Results are provided in Table 5. For each image, we zero out the activations of its class-related concepts and report the Intervention R (Int. R) metric. As shown in Table 5, this intervention reduces accuracy by approximately 20% on average, underscoring the concepts importance to the CBM.
Model Orig. CBM Int. (R)$\downarrow$ Int. (K)$\uparrow$
BeiT-B/16 54.29 41.43 (-12.86) 58.57 (+4.28)
[…]
DiNOv2 52.86 43.57 (-9.29) 59.29 (+6.43)
BeiT-L/16 52.86 44.29 (-8.57) 58.57 (+5.71)
Table 4: CBM Interventions on the standard Waterbirds dataset
Model Orig. CBM Int. (R)$\downarrow$
ResNet50 96.80 76.00 (-20.8)
ResNet101 97.00 76.80 (-20.2)
DINOv2-B 98.60 78.40 (-20.2)
BeiT-B/16 98.60 78.80 (-19.8)
[…]
Table 5: CBM Interventions in a multi-class setup
Appendix F Performance on Additional Models
Model Top-1 Orig. $\Delta$
ConvNeXt-Tiny 82.19 82.52 --0.33
ViT-B/32 75.40 75.91 --0.51
Swin-Small 82.63 83.20 --0.57
Swinv2-Tiny 81.44 82.07 --0.63
CvT-21 80.45 81.27 --0.82
Swinv2-Small 83.32 83.71 --0.39
[…]
Appendix G Implementation Details
For the text encoder, we use the all-MiniLM-L12-v1222https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library [58]. This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs. It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization [61], a GELU activation function [62] and Dropout [63] with a drop probability of 0.5. The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function. The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder). We train the MLP projector with a batch size of 256 using the ADAM optimizer [64] with a learning rate of 1e-4 that decays using a cosine schedule [65] over the total number of epochs. We follow the original image sizes that the classifier was trained on.
Appendix H Process of Decoding Visual Features
We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$. A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$. To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning [20], which prepends learnable tokens in the embedding space. We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly. Our method builds upon the work of [3].
With the updated prefixes, $G$ is run again, and the most probable token is selected as the next word. This process is repeated for a predefined number of timesteps (up to the desired sentence length) or until the <.><.>< . > token is generated. At the end of each iteration, a full sentence is generated. We conduct this process for 20 iterations, generating 20 sentences in total. The final output is chosen as the sentence with the highest similarity to the visual features $\tilde{f}$.
We also add the fluency loss from [3] as well as other token processing operations. We refer readers to [3] for more information. We use the small GPT-2 of 124M parameters as $G$. We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
Figure 2: The process used to decode visual features of an image. The process is shown for the first timestep $ts=1$ with a hard prompt set as “an image of a".
Appendix I Using only class names
We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP. In practice, we can go beyond class names by using resources like a class hierarchy from WordNet [66] (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL [67] or VCVD[39]. However, this approach would be considered as “cheating. The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes. Explicitly providing additional information would not replicate the classifier faithfully since it would force the classifier to focus on predefined features or those we intend it to learn. Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation. For instance, if class descriptions were used in the training, the concepts in CBMs would align with those specified in the training prompts. For these reasons, we refrain from using any other additional information than the class names. We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
Appendix J Compositional Captioning
The prompt we used for compositional captioning is as follows:
J.1 Domain-Specific Compositional Captioning
Next, we explore alternative concept sets in Compositional Captioning. In the main manuscript, we reported results using the 20,000 most common English words as our concept set. Since the LLM remains fixed and functions as a composer, integrating detected concepts and verbs grounded in the image into a caption, we can seamlessly substitute the concept set with any domain-specific concept set alternative. This allows for the generation of captions (here, decoded visual features) tailored to a specific domain. Here, we maintain the same set of verbs but explore the use of concepts specific to the ImageNet dataset. Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset. We use the ImageNet-specific concept set from [15] and report zero-shot captioning performance in Table 7. As shown, we observe a decrease in all metrics. This shows that our method can readily decode visual features into text for any domain. Finally, also note that we can control the style of the generations by simply prompting the LLM to compose the concepts and verbs in a specific style (e.g., humorous, positive, negative).
@@ Appendix I Using only class names
Appendix K Global Class-Wise Qualitative Examples
@@
Appendix L Waterbirds/Landbirds Dataset
There are two challenges associated with using the Waterbirds-100 dataset [59] in our work. First, our method cannot transform a classifier trained on this dataset because learning meaningful text representations from merely two labels (waterbird and landbird) with our MLP is not feasible. Second, we cannot evaluate our ImageNet-trained models on this dataset due to its highly artificial nature, which results in numerous severe out-of-distribution samples for classifiers trained on ImageNet. Therefore, we manually curated our own waterbird/landbird dataset using ImageNet validation images. Specifically, for the waterbird images, we consider classes of birds from ImageNet that are at least 90% of the time found in water backgrounds— a statistic verified by manually inspecting 100 random training images of those birds. For those birds, we then select their images in the ImageNet validation set that appear on land backgrounds. We perform a similar procedure for the landbird class. For the landbird class, we encountered a problem. Although many ImageNet bird classes nearly always feature land backgrounds in their training images, we were unable to find a substantial number of validation images depicting these birds against water backgrounds (e.g., we could not find any image of the bird robin with water background in the ImageNet validation set). In order to solve this issue, we utilized the Stable Diffusion 2.1 [68] text-to-image generative model to generate images of those birds on water backgrounds. We ensured that the generated images of those birds have the correct physical and distinctive features of the bird. For all images, we always ensure that the background (water/land) is clearly visible. This leads us to a validation dataset of 140 images (70 images for each class). Specifically, for the waterbird class, all 70 images coming from the ImageNet validation set. For the landbirds class, 15 images come from the ImageNet validation set, and 55 are generated with Stable Diffusion.
Appendix M Multi-class CBM Intervention Classes and Concepts

### v3

@@ Appendix I Other Concept Sets
Table 10: Zero-Shot Compositional Captioning Performance
While the results on CiDEr and SPICE are incremental compared to the results in Table 3, the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis that the low scores of B4 and M were attributed to the specific caption style and structure of the respective dataset (COCO).
K.1 Domain-Specific Compositional Captioning
Since the LLM we used for compositional image captioning is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains. We explore alternative concept sets in Compositional Captioning. In the main manuscript, we reported results using the 20,000 most common English words as our concept set. Since the LLM remains fixed and functions as a composer, integrating detected concepts and verbs grounded in the image into a caption, we can seamlessly substitute the concept set with any domain-specific concept set alternative. This allows for the generation of captions tailored to a specific domain. Here, we maintain the same set of verbs but explore the use of concepts specific to the ImageNet dataset. Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset. We use the ImageNet-specific concept set from (oikarinen2023labelfree) and report zero-shot captioning performance in Table 11. As shown, we observe a decrease in all metrics. This shows that our method can readily produce captions for any domain. Finally, also note that we can control the style of the generations by simply prompting the LLM to compose the concepts and verbs in a specific style (e.g., humorous, positive, negative).
@@
Appendix L Performance on Additional Models
Model Top-1 Orig. $\Delta$
ConvNeXt-Tiny 82.19 82.52 -0.33
ViT-B/32 75.40 75.91 -0.51
Swin-Small 82.63 83.20 -0.57
Swinv2-Tiny 81.44 82.07 -0.63
CvT-21 80.45 81.27 -0.82
Swinv2-Small 83.32 83.71 -0.39
[…]
Appendix M Process of Zero-Shot Image Captioning
We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$. A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$. To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning (Li2021PrefixTuningOC), which prepends learnable tokens in the embedding space. We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly. Our method builds upon the work of (Tewel2021ZeroCapZI).
With the updated prefixes, $G$ is run again, and the most probable token is selected as the next word. This process is repeated for a predefined number of timesteps (up to the desired sentence length) or until the <.><.> token is generated. At the end of each iteration, a full sentence is generated. We conduct this process for 20 iterations, generating 20 sentences in total. The final output is chosen as the sentence with the highest similarity to the visual features $\tilde{f}$.
We also add the fluency loss from (Tewel2021ZeroCapZI) as well as other token processing operations. We refer readers to (Tewel2021ZeroCapZI) for more information. We use the small GPT-2 of 124M parameters as $G$. We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
Figure 2: The process used to generate zero-shot captions using any pretrained language decoder (e.g., GPT-2). The process is shown for the first timestep ($ts=1$) and first iteration ($j=1$) with a hard prompt set as “an image of a". We apply prefix tuning while keeping the language decoder frozen, generating text that maximizes the similarity with the visual features.
Appendix N Concept Filtering
One of our concept filtering procedures requires us to know the terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name. In order to obtain this information, we used an LLM (gpt-4o-mini) with the following prompt:
Provide your answer to below as single-word comma separated. If you need to provide a term composed of 2 words, then separate each into a single word. For each class, provide its synonyms and closely related names (e.g., other species of the category, its superclasses such as bird, fish, dog, cat, animal...etc). Here is an example. The ImageNet class is: tench, tinca tinca. Answer: fish,animal,cyprinid,carp,vertebrate.
Appendix O Using only class names
We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP. In practice, we can go beyond class names by using resources like a class hierarchy from WordNet (Lin1998WordNetAE) (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL (Pratt2022WhatDA) or VCVD(Menon2022VisualCV). However, this approach would be considered as “cheating. The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes. Explicitly providing additional information would not replicate the classifier faithfully since it would force the classifier to focus on predefined features or those we intend it to learn. Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation. For instance, if class descriptions were used in the training, the concepts in CBMs would align with those specified in the training prompts. For these reasons, we refrain from using any other additional information than the class names. We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
Appendix P Global Class-Wise Qualitative Examples
@@
Appendix Q Multi-class CBM Intervention Classes and Concepts